# Skill-Enhanced Reinforcement Learning Acceleration from Demonstrations

### v1

@@ Abstract
Learning from Demonstration (LfD) aims to facilitate rapid Reinforcement Learning (RL) by leveraging expert demonstrations to pre-train the RL agent. However, the limited availability of expert demonstration data often hinders its ability to effectively aid downstream RL learning. To address this problem, we propose a novel two-stage method dubbed as Skill-enhanced Reinforcement Learning Acceleration (SeRLA). SeRLA introduces a skill-level adversarial Positive-Unlabeled (PU) learning model to extract useful skill prior knowledge by enabling learning from both limited expert data and general low-cost demonstration data in the offline prior learning stage. Subsequently, it deploys a skill-based soft actor-critic algorithm to leverage this acquired prior knowledge in the downstream online RL stage for efficient training of a skill policy network. Moreover, we develop a simple skill-level data enhancement technique to further alleviate data sparsity and improve both skill prior learning and downstream skill policy training. Our experimental results on multiple standard RL environments show the proposed SeRLA method achieves state-of-the-art performance on accelerating reinforcement learning on downstream tasks, especially in the early learning phase.
@@ 1 Introduction
Despite the wide applicability of reinforcement Learning (RL) across robotics [1], video games [2, 3], and large language models [4, 5], a conventional deep RL agent often requires numerous iterative interactions with the environment to learn a useful policy by maximizing the expected discounted cumulative reward [6], resulting in prolonged training periods and limited computational efficiency that can be particularly problematic for large input scales. To overcome this problem, Learning from Demonstration (LfD), also known as imitation learning (IL), has been investigated to accelerate RL. In LfD, the agent is pre-trained on a small offline demonstration dataset provided by human experts [7, 8] to acquire knowledge and learn behaviors that can be executed in the environment, which can then be leveraged to accelerate the online learning process of the downstream RL task with fewer interactions with the environment. Due to the limited availability of expert demonstration data, some recent studies seek to supplement the expert data with a large task-agnostic demonstration dataset collected inexpensively using methods such as autonomous exploration [9, 10] or human-tele operation [11, 12].
@@
As an RL technique that learns reusable skills from a given expert behavior [13, 14] or interacting with online environments [15, 16, 14, 9] to guide the RL training, skill-based RL shows great potential for LfD. Recently, researchers have introduced skill-based RL to LfD by learning reusable skill behaviors from the demonstration data and deploying them for the downstream tasks [11, 17, 18, 19, 20]. However, these previous studies either focus solely on learning from the expert dataset or treat the general demonstration data as negative examples, impeding the effective utilization of the low-cost demonstration data that are readily available and can still contain valuable fragmented skills.
@@
In this paper, we propose a novel SeRLA method, which stands for Skill-enhanced Reinforcement Learning Acceleration from demonstrations, to address the problem of learning from heterogeneous demonstration data and accelerating the downstream RL with the learned knowledge. SeRLA accelerates RL by pursuing skill-level learning in two stages with three coherent components: a skill-level adversarial Positive-Unlabeled (PU) learning model, a skill-based policy learning algorithm, and a skill-level data enhancement technique. In the offline skill prior training stage, the skill-level adversarial PU learning model induces useful skill prior knowledge by exploiting the general and task-agnostic demonstration data as unlabeled examples in addition to the positive expert data, instead of simply differentiating them. This strategy facilitates improved utilization of the extensive low-cost demonstration data and help alleviate the scarcity of the expert data. In the online downstream RL policy training stage, a skill-based soft actor-critic algorithm is deployed to integrate skills learned in the offline stage and accelerate skill policy learning. Moreover, a simple but novel Skill-level Data Enhancement (SDE) technique is introduced to improve the robustness of skill learning and adaptation at both stages. We conduct experiments on four standard RL environments by comparing the proposed SeRLA with the state-of-the-art skill-based imitation learning methods: SPiRL [17] and model-based SkiMo [18].
The main contributions of this paper can be summarized as follows:
@@
  • This is the first work that conducts skill-level Positive-Unlabeled Learning for LfD/IL. The proposed SeRLA takes the low-cost general demonstration data as unlabeled examples to statistically support skill learning from the limited positive examples (i.e., the expert demonstration data) through skill-level adversarial PU learning.
@@ 2 Related Works
RL from Demonstrations
@@
Learning from Demonstration (LfD) or imitation learning (IL), unlike offline RL which learns optimal policies solely from offline data, focuses on accelerating downstream RL tasks by pre-training the RL agent on a small set of expert demonstrations without reward signals [7, 8]. In addition to the limited expert demonstrations, large task-agnostic demonstration datasets can also be collected inexpensively [9, 10, 11, 12] from the environment for extracting potential learnable behaviors through LfD. There are three major methods to solve LfD/IL problems: Behavior Cloning (BC) [7], Inverse Reinforcement Learning (IRL) [21], and Generative Adversarial Imitation Learning (GAIL) [22]. BC learns a direct mapping from observations to actions through supervised learning but struggles with generalization and distribution shifts [23, 24]. IRL infers reward functions from demonstrations and trains the agent using standard RL algorithms, but it is computationally expensive and relies on the reward model’s effectiveness [25]. GAIL uses a generative adversarial network where a discriminator distinguishes between agent and expert behaviors, achieving strong performance despite needing many interactions with the environment [26].
Skill-Based RL
@@
Skill-based RL methods extract reusable skills as abstracted long-horizon behavior sequences of actions [13, 15, 17, 27, 14, 9, 11, 16], where the skills are predefined by experts or extracted from datasets. In recent works, skill-Prior RL (SPiRL) accelerates downstream RL tasks using learned skill priors from offline demonstration data [17]. Skill-based Learning with Demonstration (SkiLD) regularizes policy learning in downstream tasks using a skill posterior [28]. Few-shot Imitation with Skill Transition Models (FIST) learns skills from few-shot demonstration data and generalizes to unseen tasks [19]. Adaptive Skill Prior for RL (ASPiRe) adaptively learns distinct skill priors from different datasets [20]. Skill-based Model-based RL (SkiMo) applies planning on downstream tasks using a skill dynamics model [18].
Positive-Unlabeled Learning
@@
Unlike standard binary classifiers that use positive and negative examples, PU learning utilizes only positive and unlabeled data [29], where unlabeled examples can belong to either class. PU learning has proven valuable in applications like medical diagnosis [30] and knowledge base construction [31, 32]. Prior work has focused on loss functions and optimizers [33, 34]. Kiryo et al. [35] extended PU learning to deep neural networks, and Xu and Denil [36] applied it to imitation reward learning by replacing the adversarial loss in GAIL [22] with PU loss, creating a PU-based reward function for RL agents. By contrast, our work delivers a skill-level two-stage training framework that learns skill priors simultaneously from both limited expert data and low-cost general demonstration data via a PU learning model to accelerate the downstream RL tasks.
@@ 3 Problem Setting
Reinforcement Learning from Demonstrations (LfD) aims to accelerate the online downstream RL procedure by leveraging offline demonstration datasets. Specifically, we assume LfD has access to two demonstration datasets: a limited expert dataset $D_{\pi_{e}}$ and a low-cost general demonstration dataset $D_{\pi}$. The expert dataset is a task-specific small offline dataset that contains expert demonstration trajectories (state-action sequences) $D_{\pi_{e}}=\{s_{0},a_{0},\cdots,s_{t},a_{t},\cdots\}$, which are generated by human experts or fully trained RL agents. The general demonstration dataset is a much larger task-agnostic offline dataset that consists of randomly collected trajectories $D_{\pi}=\{s_{0},a_{0},\cdots,s_{t},a_{t},\cdots\}$. The action sequences contained in $D_{\pi_{e}}$ and $D_{\pi}$ can be denoted as $A_{\pi_{e}}$ and $A_{\pi}$, respectively. While the general demonstration dataset does not provide as much useful information as the expert dataset, it may still contain short-horizon behaviors that, if properly extracted, can guide the RL agent to behave with feasible actions and propel the policy learning.
The downstream RL task is a standard reinforcement learning problem that can be represented as a Markov Decision Process (MDP) $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, as described in [6]. In this MDP, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is the transition dynamics $p(s_{t+1}|s_{t},a_{t})$, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, and $\gamma\in(0,1)$ is the discount factor. The goal is to learn an optimal policy $\pi^{\star}:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted cumulative reward (return): $\pi^{\star}=\arg\max_{\pi}\;J_{r}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}]$. The goal of this work is to learn useful skill prior from the offline demonstration datasets and then deploys such skill-level knowledge to facilitate fast policy training for the downstream RL task.
@@ 4 Method
The proposed SeRLA method has two stages: the offline skill prior training stage with skill-level adversarial PU learning and the online skill-based downstream policy training. The skill prior training induces useful high-level skill knowledge in form of skill prior from the given demonstration datasets ($D_{\pi_{e}}$ and $D_{\pi}$), which is then used to accelerate the downstream skill-based policy training through a Skill-based Soft Actor-Critic (SSAC) algorithm. Moreover, a simple skill-level data enhancement technique is further devised for the two training stages to improve the overall performance. Below, we elaborate on these approach components.
4.1 Skill Prior Training with PU Learning
In the skill prior training stage, we build a regularized deep autoencoder model with skill-level adversarial PU learning to learn a conditional skill prior distribution function $q_{\psi}(z_{t}|s_{t})$ from the trajectories provided in the two demonstration datasets, where latent variables $\{z_{t}\}$ are used to capture the high level representations of skills, each of which can be interpreted as an action sequence. The model includes a skill encoder network $q_{\mu}(\cdot)$, a skill decoder network $p_{\nu}(\cdot)$, a skill prior network $q_{\psi}(\cdot)$, and a discriminator $\mathcal{D}_{\zeta}$. The first three components can be learned from the expert data $D_{\pi_{e}}$ within a conventional autoencoder framework, while the discriminator is innovatively deployed to incorporate the general demonstration data $D_{\pi}$ into the skill prior training process and alleviate skill data sparsity via adversarial PU learning.
@@ 4.1 Skill Prior Training with PU Learning
Under a deep autoencoder framework [17], the skill encoder $q_{\mu}(z_{t}|{\bf a}_{t})$ takes an action sequence ${\bf a}_{t}=\{a_{t},...,a_{t+H-1}\}$ with length $H$ as input and maps it to a latent skill embedding $z_{t}$. Conversely, the skill decoder $p_{\nu}(\hat{\bf a}_{t}|z_{t})$ reconstructs an action sequence $\hat{\textbf{a}}_{t}$ from the given skill $z_{t}$. The autoencoder can be learned by minimizing a reconstruction loss on the expert action sequences $A_{\pi_{e}}$:
@@
where $\mathcal{L}_{KL}(\cdot,\cdot)$ denotes the Kullback Leibler (KL) divergence function. Moreover, a standard Gaussian distribution prior $p(z_{t})$ $=$ $\mathcal{N}(0,1)$ can be deployed to regularize the skill embedding space with the following regularization loss:
@@ 4.1.1 Skill-level Adversarial PU Learning
Different from the expert data, the randomly collected large demonstration data $D_{\pi}$ can present a great number of short-horizon behaviors, some of which can be meaningful while many others can be spontaneous or arbitrary. Hence it is not suitable to directly deploy $D_{\pi}$ in the autoencoder model above in the same way as the expert data. To effectively filter out the noisy behaviors and exploit the useful ones from $D_{\pi}$, we deploy a PU learning scheme to perform skill learning simultaneously from both the small expert data $D_{\pi_{e}}$ and the large general demonstration data $D_{\pi}$. Specifically, we treat the skills (capturing the behaviors) from the expert data, $Z_{e}=q_{\mu}(A_{\pi_{e}})$, as positive examples (i.e., useful skills), and treat skills from the general demonstration data, $Z=q_{\mu}(A_{\pi})$, as unlabeled examples that can be either positive or negative. Then we propose to learn a binary probabilistic discriminator $\mathcal{D}_{\zeta}$ from the positive and unlabeled skill examples by adapting a standard non-negative PU risk function derived in the literature [35] into the skill-level learning:
${L}_{\mathcal{D}_{\zeta}}^{pu}(q_{\mu}(A_{\pi_{e}}),q_{\mu}(A_{\pi}))=\lambda{L}_{\mathcal{D}_{\zeta}}^{1}(q_{\mu}(A_{\pi_{e}}))+\max(-\xi,{L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi}))-\lambda{L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi_{e}})))$ (4)
where $\lambda>0$ and $\xi\geq 0$ are hyperparameters. Here the true positive risk ${L}_{\mathcal{D}_{\zeta}}^{1}(q_{\mu}(A_{\pi_{e}}))$ is calculated on positive skill examples $Z_{e}$, while the true negative risk is calculated on both positive and unlabeled data ($Z_{e}$ and $Z$) using two terms, ${L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi_{e}}))$ and ${L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi}))$. These risk terms are defined in terms of the discriminator $\mathcal{D}_{\zeta}$ as follows:
@@
Output: Trained skill prior network $q_{\psi}(z_{t}|s_{t})$, and decoder $p_{\nu}(a_{t:t+H-1}|z_{t})$
@@
3:         Sample $s_{t}$ and sequence $a_{t:t+H-1}$ from $D_{\pi_{e}}$
4:         Sample action sequence $a_{t^{\prime}:t^{\prime}+H-1}^{\prime}$ from $D_{\pi}$
@@
$L(\mu,\nu,\psi)=L_{rec}(\nu,\mu)+L_{prior}(\psi,\mu)+\beta L_{reg}(\mu)-\rho\min\nolimits_{\zeta}{L}_{\mathcal{D}_{\zeta}}^{pu}(q_{\mu}(A_{\pi_{e}}),q_{\mu}(A_{\pi}))$ (8)
where the reconstruction loss $L_{rec}$ enforces consistency between the skill embedding $z_{t}$ and the action sequence ${\bf a}_{t}$; the prior training loss $L_{prior}$ ensures that the generated skill is consistent with the current state and action sequence; $L_{reg}$ regularizes the skill embedding space; and the PU loss ${L}_{D_{\zeta}}^{pu}$ is used to effectively incorporate large random demonstration data into skill learning in an adversarial manner. The joint training of all these components is expected to effectively and proficiently learn valuable skill knowledge from the unified heterogeneous offline demonstration datasets.

### v2

# Skill-Enhanced Reinforcement Learning Acceleration from Heterogeneous Demonstrations

@@ Abstract
Learning from Demonstration (LfD) is a well-established problem in Reinforcement Learning (RL), which aims to facilitate rapid RL by leveraging expert demonstrations to pre-train the RL agent. However, the limited availability of expert demonstration data often hinders its ability to effectively aid downstream RL learning. To address this problem, we propose a novel two-stage method dubbed as Skill-enhanced Reinforcement Learning Acceleration (SeRLA). SeRLA introduces a skill-level adversarial Positive-Unlabeled (PU) learning model that extracts useful skill prior knowledge by learning from both expert demonstrations and general low-cost demonstrations in the offline prior learning stage. Building on this, it employs a skill-based soft actor-critic algorithm to leverage the acquired priors for efficient training of a skill policy network in the downstream online RL stage. In addition, we propose a simple skill-level data enhancement technique to mitigate data sparsity and further improve both skill prior learning and skill policy training. Experiments across multiple standard RL benchmarks demonstrate that SeRLA achieves state-of-the-art performance in accelerating reinforcement learning on downstream tasks, particularly in the early training phase.
[…]
4294
@@ 1 Introduction
Despite the wide applicability of Reinforcement Learning (RL) across robotics [21], video games [39, 4], and large language models [37, 29], a conventional deep RL agent often requires numerous iterative interactions with the environment to learn a useful policy by maximizing the expected discounted cumulative reward [38], resulting in prolonged training periods and limited computational efficiency. This challenge becomes more pronounced in complex environments, where exploration is both costly and time-consuming. To overcome this problem, Learning from Demonstration (LfD), also known as imitation learning (IL), has been investigated to accelerate RL. In LfD, the agent is pre-trained on a small offline demonstration dataset provided by human experts [2, 5] to acquire knowledge and learn behaviors that can be executed in the environment, which can then be leveraged to accelerate the online learning process of downstream RL tasks with fewer environment interactions. Due to the limited availability of expert demonstration data, some recent studies seek to supplement the expert data with a large task-agnostic demonstration dataset collected inexpensively using methods such as autonomous exploration [18, 35] or human-teleoperation [15, 27].
@@
Skill-based RL, which acquires reusable skills—high-level behaviors composed of primitive actions—from expert demonstrations [22, 9] or environment interactions [23, 11, 9, 18] to guide reinforcement learning, shows great potential for advancing LfD. Recently, researchers have introduced skill-based RL to LfD by learning reusable skill behaviors from demonstration data and deploying them for downstream tasks [15, 31, 36, 17, 41]. However, these previous studies either focus solely on learning from expert datasets or treat general demonstration data as negative examples, thereby impeding the effective utilization of low-cost demonstrations that are widely available and may contain valuable fragmented skills.
@@
In this paper, we propose a novel SeRLA method, which stands for Skill-enhanced Reinforcement Learning Acceleration from heterogeneous demonstrations, to address the problem of learning from heterogeneous demonstration data and accelerating downstream RL with the learned knowledge. SeRLA accelerates RL by pursuing skill-level learning in two stages with three coherent components: a skill-level adversarial PU learning module, a skill-based soft actor-critic policy learning algorithm, and a skill-level data enhancement technique. In the offline skill prior training stage, the skill-level adversarial PU learning module learns useful skill prior knowledge by exploiting the general, task-agnostic demonstration data as unlabeled examples in addition to the positive expert data, instead of simply differentiating them. This strategy facilitates improved utilization of the extensive low-cost demonstration data and can help alleviate the scarcity of the expert data. In the online downstream RL policy training stage, a skill-based soft actor-critic algorithm is deployed to integrate skills learned in the offline stage and accelerate skill policy learning. Moreover, a simple but novel Skill-level Data Enhancement (SDE) technique is introduced to improve the robustness of skill learning and adaptation at both stages. We conduct experiments on four standard RL environments by comparing the proposed SeRLA with the state-of-the-art skill-based imitation learning methods: SPiRL [31] and model-based SkiMo [36]. The main contributions of this paper are summarized as follows:
@@
  • This is the first work that conducts skill-level Positive-Unlabeled Learning for LfD/IL. The proposed SeRLA uses low-cost, general demonstration data as unlabeled examples to statistically support skill learning from limited positive examples—i.e., expert demonstrations—through skill-level adversarial PU learning.
@@ 2 Related Works
2.1 Reinforcement Learning from Demonstrations
@@
Learning from Demonstration (LfD) or imitation learning (IL) aims to accelerate reinforcement learning for downstream RL tasks by pre-training the RL agent on a small expert demonstration dataset typically without reward signals [2, 5]. Unlike offline RL, which aims to learn optimal policies solely from offline data, LfD/IL focuses on accelerating downstream tasks. In addition to the limited expert demonstrations in the form of a sequence of state-action pairs $\{(s_{0},a_{0}),...,(s_{t},a_{t})\}$, large task-agnostic demonstration datasets can also be collected inexpensively [18, 35, 15, 27] from the environment for extracting potential learnable behaviors through LfD. Apart from these existing works, recent works for LfD/IL also include Behavior Cloning (BC) [2], Inverse Reinforcement Learning (IRL) [1], and Generative Adversarial Imitation Learning (GAIL) [19]. BC enables the RL agent to learn a direct mapping between observations and corresponding actions from the demonstration dataset as a supervised learning problem. This method however has limited generalization ability and suffers from distribution shifts [33, 34]. IRL infers reward functions from the demonstration data and trains the RL agent using standard RL algorithms [26]. Although IRL can transform imitation learning to a standard RL problem, it is computationally expensive and relies heavily on the effectiveness of the reward model. GAIL treats IL as a two-player zero-sum game with a generative adversarial network [14], where a discriminator is trained to distinguish the behavior between the agent policy and the expert policy learned from the demonstrations. It solves the zero-sum game using minimax optimization, yielding similar behaviors between the agent policy and the expert policy. Despite the requirement of numerous interactions with the environment, GAIL demonstrates remarkable performance on IL.
2.2 Skill-Based Reinforcement Learning
@@
As a popular approach to leveraging prior knowledge, skill-based RL methods extract reusable skills as abstracted long-horizon behavior sequences of actions [22, 23, 31, 24, 9, 18, 15, 11]. These skills can either be predefined by experts or extracted from online or offline datasets, naturally supporting LfD or IL. One recent work [31] in this line introduces Skill-Prior RL (SPiRL), a hierarchical skill-based approach, to accelerate downstream RL tasks using learned skill priors from offline demonstration data. Another work [32] proposes a Skill-based Learning with Demonstration (SkiLD) approach, which uses a skill posterior to regularize the policy learning in the downstream task. In addition, the approach of Few-shot Imitation with Skill Transition Models (FIST) [17] learns skills from few-shot demonstration data and generalizes to unseen tasks. The Adaptive Skill Prior for RL (ASPiRe), introduced by Xu et al. [41], adaptively learns distinct skill priors from different datasets with specific weights. Furthermore, Shi et al. [36] developed a Skill-based Model-based RL framework (SkiMo) that applies planning on downstream tasks by using a skill dynamics model to select the optimal skill from different learned skills. Celik et al. [6] introduced energy-based models to learn diverse skills via a mixture of experts.
2.3 Positive-Unlabeled (PU) Learning
@@
Unlike the standard binary classifier that learns from positive and negative examples, PU learning utilizes only positive and unlabeled data [3], where an unlabeled example can belong to either the positive or negative class. PU learning has demonstrated values in real-world applications with PU data such as medical diagnosis [7] and knowledge base construction [13, 43]. Prior works on PU learning focus on the loss functions and optimizers [10, 30]. A recent work by Kiryo et al. [20] builds large scale PU learning upon deep neural networks. Zhao et al. [42] proposed a novel boosting framework aimed at improving PU classifier training efficiency. More recently, Xu and Denil [40] demonstrated the utility of PU Learning in imitation reward learning by replacing the adversarial loss with a PU loss in GAIL [19] and developing a specific PU learning-based reward function to train the RL agent with expert demonstration data. By contrast, our proposed work delivers a skill-level two-stage training framework that can learn skill priors simultaneously from both limited expert data and low-cost general demonstration data via a PU learning module, while accelerating the downstream RL tasks.
Figure 1: The proposed method, SeRLA, pursues skill-level training in two stages: offline skill prior training via skill-level PU Learning and online downstream skill policy training via SSAC (Skill-based Soft Actor-Critic). Left: Skill-level PU-Learning incorporates a large random demonstration dataset $D_{\pi}$ into skill learning on a small expert dataset $D_{\pi_{e}}$ by adding a discriminator $\mathcal{D}_{\zeta}$ through an adversarial PU loss. Right: The learned skill knowledge is exploited to accelerate the downstream online RL task through a model-free SSAC algorithm, which utilizes the prior skill knowledge through behavior cloning. Moreover, Skill-level Data Enhancement (SDE) is proposed to further alleviate data sparsity and improve learning robustness by employing skill augmentation in both the offline prior learning stage and the online downstream RL training stage.
@@ 3 Problem Setting
Reinforcement Learning from Demonstrations (LfD) aims to accelerate the online downstream RL procedure by leveraging offline demonstration datasets. Specifically, we assume LfD has access to two demonstration datasets: a limited expert dataset $D_{\pi_{e}}$ and a low-cost general demonstration dataset $D_{\pi}$. The expert dataset is a task-specific small offline dataset that contains expert demonstration trajectories (state-action sequences) $D_{\pi_{e}}=\{s_{0},a_{0},\cdots,s_{t},a_{t},\cdots\}$, which are generated by human experts or fully trained RL agents. The general demonstration dataset is a much larger task-agnostic offline dataset that consists of randomly collected trajectories $D_{\pi}=\{s_{0},a_{0},\cdots,s_{t},a_{t},\cdots\}$. The action sequences contained in $D_{\pi_{e}}$ and $D_{\pi}$ can be denoted as $A_{\pi_{e}}$ and $A_{\pi}$, respectively. While the general demonstration dataset does not provide as much useful information as the expert dataset, it may still contain short-horizon behaviors that, if properly extracted, can guide the RL agent to behave with feasible actions and propel policy learning.
The downstream RL task is a standard reinforcement learning problem that can be represented as a Markov Decision Process (MDP) $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, as described in [38]. In this MDP, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is the transition dynamics $p(s_{t+1}|s_{t},a_{t})$, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, and $\gamma\in(0,1)$ is the discount factor. The goal is to learn an optimal policy $\pi^{\star}:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted cumulative reward (return): $\pi^{\star}=\arg\max_{\pi}\;J_{r}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}]$. The goal of this work is to learn useful skill prior from the offline heterogeneous demonstration datasets and deploy such skill-level knowledge to facilitate fast policy training for the downstream RL task.
@@ 4 Method
The main architecture of the proposed SeRLA method is illustrated in Figure 1. It has two stages: the offline skill prior training stage with skill-level adversarial PU learning and the online skill-based downstream policy training. The skill prior training induces useful high-level skill knowledge in form of skill priors from the given heterogeneous demonstration datasets ($D_{\pi_{e}}$ and $D_{\pi}$), which is then used to accelerate the downstream skill-based policy training through a Skill-based Soft Actor-Critic (SSAC) algorithm. Moreover, a simple skill-level data enhancement technique is further devised for the two training stages to improve the overall performance. Below, we elaborate on these approach components.
4.1 Skill Prior Training with Adversarial PU Learning
In the skill prior training stage, we build a regularized deep autoencoder module with skill-level adversarial PU learning to learn a conditional skill prior distribution function $q_{\psi}(z_{t}|s_{t})$ from the trajectories provided in the two demonstration datasets, where latent variables $\{z_{t}\}$ are used to capture the high level representations of skills, each of which can be interpreted as an action sequence. The model includes a skill encoder network $q_{\mu}(\cdot)$, a skill decoder network $p_{\nu}(\cdot)$, a skill prior network $q_{\psi}(\cdot)$, and a discriminator $\mathcal{D}_{\zeta}$. The first three components can be learned from the expert data $D_{\pi_{e}}$ within a conventional autoencoder framework, while the discriminator is innovatively deployed to incorporate the general demonstration data $D_{\pi}$ and alleviate skill data sparsity via adversarial PU learning.
@@ 4.1 Skill Prior Training with Adversarial PU Learning
Under a deep autoencoder framework [31], the skill encoder $q_{\mu}(z_{t}|{\bf a}_{t})$ takes an action sequence ${\bf a}_{t}=\{a_{t},...,a_{t+H-1}\}$ with length $H$ as input and maps it to a latent skill embedding $z_{t}$. Conversely, the skill decoder $p_{\nu}(\hat{\bf a}_{t}|z_{t})$ reconstructs an action sequence $\hat{\textbf{a}}_{t}$ from a given skill $z_{t}$. The autoencoder can be learned by minimizing a reconstruction loss on the observed action sequences $A_{\pi_{e}}$ from the expert demonstrations $D_{\pi_{e}}$:
@@
where $\mathcal{L}_{KL}(\cdot,\cdot)$ denotes the Kullback Leibler (KL) divergence function. Moreover, a standard Gaussian distribution prior $p(z_{t})$ == $\mathcal{N}(0,1)$ can be deployed to regularize the skill embedding space with the following regularization loss:
@@ 4.1.1 Skill-level Adversarial PU Learning
Different from the expert data, the randomly collected low-cost, large demonstration dataset $D_{\pi}$ can present a great number of short-horizon behaviors, some of which can be meaningful while many others can be spontaneous or arbitrary. Hence it is not suitable to directly deploy $D_{\pi}$ in the autoencoder model above in the same way as the expert data $D_{\pi_{e}}$. To effectively filter out the noisy behaviors and exploit the useful ones from $D_{\pi}$, we deploy a PU learning scheme to perform skill learning simultaneously from both the small expert data $D_{\pi_{e}}$ and the large general demonstration data $D_{\pi}$. PU learning is a variant of supervised learning that learns a binary classifier from positive and unlabeled data [3]. By modeling the unlabeled data instead of assuming it is entirely negative, PU learning can exploit additional information and reduce bias compared to naive methods. Specifically, we treat the skills (capturing the behaviors) from the expert data, $Z_{e}=q_{\mu}(A_{\pi_{e}})$, as positive examples (i.e., useful skills), and treat skills from the general demonstration data, $Z=q_{\mu}(A_{\pi})$, as unlabeled examples that can be either positive or negative. Then we propose to learn a binary probabilistic discriminator $\mathcal{D}_{\zeta}$ from the positive and unlabeled skill examples by adapting a standard non-negative PU risk function derived in the literature [20] into the skill-level learning:
${L}_{\mathcal{D}_{\zeta}}^{pu}(q_{\mu}(A_{\pi_{e}}),$ $q_{\mu}(A_{\pi}))=\lambda{L}_{\mathcal{D}_{\zeta}}^{1}(q_{\mu}(A_{\pi_{e}}))+$
$\max(-\xi,{L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi}))-\lambda{L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi_{e}})))$ (4)
where $\lambda>0$ and $\xi\geq 0$ are hyperparameters. Here the true positive risk ${L}_{\mathcal{D}_{\zeta}}^{1}(q_{\mu}(A_{\pi_{e}}))$ is calculated on positive skill examples $Z_{e}=q_{\mu}(A_{\pi_{e}})$, while the true negative risk is calculated on both positive and unlabeled data ($Z_{e}$ and $Z$) using two terms, ${L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi_{e}}))$ and ${L}_{\mathcal{D}_{\zeta}}^{0}(q_{\mu}(A_{\pi}))$. These risk terms are defined in terms of the discriminator $\mathcal{D}_{\zeta}$ as follows:
@@
Output: Trained skill prior $q_{\psi}(z_{t}|s_{t})$, and decoder $p_{\nu}(a_{t:t+H-1}|z_{t})$
@@
3:   Sample $s_{t}$ and sequence ${\bf a}_{t}=a_{t:t+H-1}$ from $D_{\pi_{e}}$
4:   Sample action sequence ${\bf a}_{t}^{\prime}=a_{t^{\prime}:t^{\prime}\!+\!H\!-\!1}^{\prime}$ from $D_{\pi}$ (or $A_{\pi}$)
@@
$L(\mu,\nu,\psi)=$ $L_{rec}(\nu,\mu)+L_{prior}(\psi,\mu)+\beta L_{reg}(\mu)$
$-\rho\min\nolimits_{\zeta}{L}_{\mathcal{D}_{\zeta}}^{pu}(q_{\mu}(A_{\pi_{e}}),q_{\mu}(A_{\pi}))$ (8)
where $\beta$ and $\rho$ are tradeoff hyperparameters; the reconstruction loss $L_{rec}$ enforces consistency between the skill embedding $z_{t}$ and the action sequence ${\bf a}_{t}$; the prior training loss $L_{prior}$ ensures that the generated skill is consistent with the current state and action sequence; $L_{reg}$ regularizes the skill embedding space; and the PU loss ${L}_{D_{\zeta}}^{pu}$ is used to effectively incorporate large random demonstration data into skill learning in an adversarial manner. The joint training of all these components is expected to effectively and proficiently learn valuable skill knowledge from the unified heterogeneous offline demonstration datasets.